# Particle Filtering for a Class of State-Space Models with Low and Degenerate Observational Noise

### v1

@@ 2.3 Noise Parametrization and Degenerate Noise
Define the manifold $\widetilde{\mathsf{M}}_{n}=\{(x_{n},\epsilon_{n})\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}:y_{n}-H_{n}(x_{n},\epsilon_{n})=0\}$. We set $c_{n}(x_{n},\epsilon_{n})=y_{n}-H_{n}(x_{n},\epsilon_{n})$. We shall suppose that $\mathbb{R}^{d_{x}}$ is equipped with a metric tensor with a fixed positive definite matrix representation $M$. Then it is simple to show that the filter density of $x_{n},\epsilon_{n}$ w.r.t. the Riemannian measure on $\widetilde{\mathsf{M}}_{n}$ denote $\mu^{M}_{\mathsf{M}_{n}}$ is as follows (see e.g. [21]). We set
$\widetilde{G}_{n}(x_{n},\epsilon_{n})=\textrm{det}(\partial c_{n}(x_{n},\epsilon_{n})M^{-1}\partial c_{n}(x_{n},\epsilon_{n})^{\top})^{-1/2}$
@@
where $\partial c_{n}$ is the Jacobian matrix and det is the determinant. For time 1 we have
@@
$\widetilde{\pi}_{1}(x_{1},\epsilon_{1})=\frac{\widetilde{G}_{1}(x_{1},\epsilon_{1})f(x_{1}|x_{0})p_{1}(\epsilon_{1})}{\int_{\widetilde{\mathsf{M}}_{1}}\widetilde{G}_{1}(x_{1},\epsilon_{1})f(x_{1}|x_{0})p_{1}(\epsilon_{1})\mu^{M}_{\widetilde{\mathsf{M}}_{1}}(d(x_{n},\epsilon_{n}))}.$
and at any subsequent time $n\geq 2$:
$\widetilde{\pi}_{n}(x_{n},\epsilon_{n})=\frac{\widetilde{G}_{n}(x_{n},\epsilon_{n})p_{n}(\epsilon_{n})\int_{\mathsf{M}_{n-1}}f(x_{n}|x_{n-1})\widetilde{\pi}_{n}(x_{n-1},\epsilon_{n-1})\mu^{M}_{\widetilde{\mathsf{M}}_{n-1}}(d(x_{n-1},\epsilon_{n-1}))}{\int_{\widetilde{\mathsf{M}}_{n}}\widetilde{G}_{n}(x_{n},\epsilon_{n})p(\epsilon_{n})\int_{\widetilde{\mathsf{M}}_{n-1}}f(x_{n}|x_{n-1})\widetilde{\pi}_{n}(x_{n-1},\epsilon_{n-1})\mu^{M}_{\widetilde{\mathsf{M}}_{n-1}}(d(x_{n-1},\epsilon_{n-1}))\mu^{M}_{\widetilde{\mathsf{M}}_{n}}(d(x_{n},\epsilon_{n}))}.$
@@
Here $p_{n}(\epsilon_{n})$ is the Lebesgue density of $\epsilon_{n}$. We remark that one can work with the noise parameters of the hidden state as well. However, as we focus on observational noise, this representation will suffice. As we will establish in the next Section when the noise in the observation equation is low, one can consider numerical algorithms under the noise representation and ultimately, as the ‘noise disappears’ recover an algorithm associated to the degenerate noise filtering problem, the latter of which we now explain
@@ 2.3.1 Degenerate Noise
The approach here is easily adapted to the case of degenerate observation noise as we now explain. Writing the SSM as
$Y_{n}$ $=H_{n}^{\star}(X_{n})$
$X_{n}$ $=F_{n}(X_{n-1},\nu_{n})$
@@
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}+d_{x}}$, $(d_{x},d_{y})\in\mathbb{N}^{2}$, $d_{y}<d_{x}$, $\nu_{n}\in\mathbb{R}^{d_{x}}$ a sequence of independent noises (independent across $n$) with positive and smooth Lebesgue density, $H_{n}^{\star}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$, $F_{n}:\mathbb{R}^{2d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ are a sequence of continuously differentiable functions and whose Jacobian is of full row rank. Note that we require $d_{y}<d_{x}$ as we aim to work with an embedded manifold.
@@
As above, define the manifold $\mathsf{M}_{n}^{\star}=\{x_{n}\in\mathbb{R}^{d_{x}}:y_{n}-H_{n}^{\star}(x_{n})=0\}$. We set $c_{n}(x_{n})=y_{n}-H_{n}^{\star}(x_{n})$. We shall suppose that $\mathbb{R}^{d_{x}}$ is equipped with a metric tensor with a fixed positive definite matrix representation $M$. We set
$G_{n}(x_{n})=\textrm{det}(\partial c_{n}(x_{n})M^{-1}\partial c_{n}(x_{n})^{\top})^{-1/2}.$
As before the filter densities w.r.t. $\mu_{M}^{\mathsf{M}_{n}^{\star}}$ can be written down as
$\pi_{1}^{\star}(x_{1})=\frac{G_{1}(x_{1})f_{1}(x_{1}|x_{0})}{\int_{\mathsf{M}_{1}^{\star}}G_{1}(x_{1})f_{1}(x_{1}|x_{0})\mu^{M}_{\mathsf{M}_{1}^{\star}}(dx_{n})}.$
and at any subsequent time $n\geq 2$:
$\pi_{n}^{\star}(x_{n})=\frac{G_{n}(x_{n})\int_{\mathsf{M}_{n-1}^{\star}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}^{\star}(x_{n-1})\mu^{M}_{\mathsf{M}_{n-1}^{\star}}(dx_{n-1})}{\int_{\mathsf{M}_{n}^{\star}}\tilde{G}_{n}(x_{n})\int_{\mathsf{M}_{n-1}^{\star}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}^{\star}(x_{n-1})\mu^{M}_{\mathsf{M}_{n-1}^{\star}}(dx_{n-1})\mu^{M}_{\mathsf{M}_{n}^{\star}}(dx_{n})}.$
3 Particular Cases
@@
In this Section we consider the case of general additive noise (Section 3.1) and derive PFs in the low and degenerate noise cases. We then examine (Section 3.2) the case where the hidden process has additive Gaussian noise along with the observation process as well and derive the optimal proposals for the low and degenerate noise models.
3.1 General Additive Noise
We consider the state-space model for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$ given:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}\\$
@@ 3.1 General Additive Noise
where, $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, $\epsilon_{n}\in\mathbb{R}^{d_{y}}$ are a sequence of independent noises with positive Lebesgue density $p_{n}$ and $\Delta\geq 0$. For the hidden dynamics, we do not make any constraints beyond the existence of a positive Lebesgue transition density which can be evaluated pointwise; it is written as $f_{n}(x_{n}|x_{n-1})$ as before. In order to write down a PF we will write the smoother (in the noise parameterization) w.r.t. Lebesgue measure as, due to the linear constraint, this is simple to do.
@@
In this article we present noise degeneracy as the case that in the observation equation, as $\Delta\downarrow 0$ there is no noise at all. However, one could consider, say $Y_{n}\in\mathbb{R}^{3}$ and an observation equation of the type
$Y_{n}=A_{n}X_{n}+\mathtt{P}(\Delta)\epsilon_{n}$
@@
In this case one could send $\Delta\downarrow 0$ which corresponds to removing the observation noise in the third observation component. This type of scenario could be dealt with in the forthcoming commentary, in the sense that one would define a filter (noise parameterization) and a corresponding algorithm in the low noise case and the case $\Delta=0$ and that there is some convergence of the low-noise case to this limiting and degenerate case. The main changes would be mainly notational and hence, for ease of exposition, we keep the simple scenario that we have started with.
3.1.1 Low Noise Case
We begin in the low noise case. Let $x_{n}^{\star}$ solve $y_{n}=A_{n}x$, $\text{ker}(A_{n})=\{x\in\mathbb{R}^{d_{x}}:A_{n}x=0\}$ and $V_{n}$ is the $d_{x}\times(d_{x}-d_{y})$ matrix whose columns are an orthonormal basis of $\text{ker}(A_{n})$. In this case we will be filtering a random variable $\widetilde{Z}_{n}\in\mathbb{R}^{d_{x}}$ and as it will prove useful below, we set $\widetilde{z}_{n}^{\top}=(z_{n}^{\top},\overline{z}_{n}^{\top})^{\top}$, with $(z_{n},\overline{z}_{n})\in\mathbb{R}^{d_{x}-d_{y}}\times\mathbb{R}^{d_{y}}$. Let $A_{n}^{\Delta}:=[A_{n},\Delta^{1/2}\mathtt{I}_{d_{y}}]$, with $\mathtt{I}_{d_{y}}$ the $d_{y}\times d_{y}$ identity matrix, be a $d_{y}\times(d_{x}+d_{y})$ matrix, define $\text{ker}(A_{n}^{\Delta}):=\{x\in\mathbb{R}^{d_{x}+d_{y}}:A_{n}^{\Delta}x=0\}$, let $V_{n}^{\Delta}$ be a $(d_{x}+d_{y})\times d_{x}$ matrix whose columns are an orthonormal basis of $\text{ker}(A_{n}^{\Delta})$ and $(\widetilde{x}_{n}^{\Delta})^{\top}=((x_{n}^{\star})^{\top},0^{\top})$ be a $(d_{x}+d_{y})$-dimensional vector that solves $y_{n}=A_{n}^{\Delta}\widetilde{x}_{n}^{\Delta}$. For any $z\in\mathbb{R}^{d_{x}}$ set
$u_{n}^{\Delta}(z)=\widetilde{x}_{n}^{\Delta}+V_{n}^{\Delta}z$
with $u_{n}^{\Delta}(z)^{\top}=(u_{n}^{\Delta}(z,x)^{\top},u_{n}^{\Delta}(z,\epsilon)^{\top})$, $u_{n}^{\Delta}(z,x)\in\mathbb{R}^{d_{x}}$, $u_{n}^{\Delta}(z,\epsilon)\in\mathbb{R}^{d_{y}}$ and we note that $y_{n}=A_{n}^{\Delta}u_{n}^{\Delta}(z)$.
Given this construction, at any time $n\geq 1$ we can write the joint smoothing density of $\widetilde{z}_{1:n}\in\mathbb{R}^{nd_{x}}$ w.r.t. $\bigotimes_{i=1}^{n}\lambda_{d_{x}}(d\widetilde{z}_{i})$ as

### v2

@@ 2.3.1 Noise Parametrization
In detail, we assume that the ambient space $\mathbb{R}^{d_{x}+d_{y}}$ is equipped with a metric tensor with a fixed positive definite matrix representation $M$. Then, it is simple to show that the filter density of $(x_{n},\epsilon_{n})$ w.r.t. the Riemannian measure on $\widetilde{\mathsf{M}}_{n}$ under $M$, denoted $\mu_{\widetilde{\mathsf{M}}_{n}}$, is as follows (see e.g. [30]). We set:
$\widetilde{G}_{n}(x_{n},\epsilon_{n})=\textrm{det}(\partial c_{n}(x_{n},\epsilon_{n})M^{-1}\partial c_{n}(x_{n},\epsilon_{n})^{\top})^{-1/2},$ (2.4)
@@
where $\partial c_{n}$ is the Jacobian matrix and $\textrm{det}(\cdot)$ is the determinant. For time 1 we have the filtering density:
$\widetilde{\pi}_{1}(x_{1},\epsilon_{1})\equiv\frac{d\widetilde{\pi}_{1}}{d\mu_{\widetilde{\mathsf{M}}_{1}}}(x_{1},\epsilon_{1})=\frac{\widetilde{G}_{1}(x_{1},\epsilon_{1})f_{1}(x_{1}|x_{0})p_{1}(\epsilon_{1})}{\int_{\widetilde{\mathsf{M}}_{1}}\widetilde{G}_{1}(x_{1},\epsilon_{1})f_{1}(x_{1}|x_{0})p_{1}(\epsilon_{1})\mu_{\widetilde{\mathsf{M}}_{1}}(d(x_{n},\epsilon_{n}))}.$
At any subsequent time $n\geq 2$, we have:
$\widetilde{\pi}_{n}(x_{n},\epsilon_{n})\equiv\frac{d\widetilde{\pi}_{n}}{d\mu_{\widetilde{\mathsf{M}}_{n}}}(x_{n},\epsilon_{n})$
$\qquad=\frac{\widetilde{G}_{n}(x_{n},\epsilon_{n})p_{n}(\epsilon_{n})\int_{\widetilde{\mathsf{M}}_{n-1}}f_{n}(x_{n}|x_{n-1})\widetilde{\pi}_{n}(x_{n-1},\epsilon_{n-1})\mu_{\widetilde{\mathsf{M}}_{n-1}}(d(x_{n-1},\epsilon_{n-1}))}{\int_{\widetilde{\mathsf{M}}_{n}}\widetilde{G}_{n}(x_{n},\epsilon_{n})p_{n}(\epsilon_{n})\int_{\widetilde{\mathsf{M}}_{n-1}}f_{n}(x_{n}|x_{n-1})\widetilde{\pi}_{n}(x_{n-1},\epsilon_{n-1})\mu_{\widetilde{\mathsf{M}}_{n-1}}(d(x_{n-1},\epsilon_{n-1}))\mu_{\widetilde{\mathsf{M}}_{n}}(d(x_{n},\epsilon_{n}))}.$
We note that under the law $d\widetilde{\pi}_{n}(x_{n},\epsilon_{n})$, $n\in\mathbb{N}$, the marginal of $x_{n}$ is as given in expressions (2.1)–(2.2) for the natural parametrization. As before one can also write a path smoothing density
@@
$\widetilde{\pi}_{n}(x_{0:n},\epsilon_{0:n})\equiv\frac{d\widetilde{\pi}_{n}}{d\mu_{\bigotimes_{k=1}^{n}\widetilde{\mathsf{M}}_{k}}}(x_{0:n},\epsilon_{0:n})\propto\prod_{k=1}^{n}\widetilde{G}_{k}(x_{k},\epsilon_{k})p_{k}(\epsilon_{k})f_{k}(x_{k}|x_{k-1}),$ (2.5)
which will form the target density of a PF.
@@
We remark that one can obtain some corresponding re-parametrized expressions by working with the noise variables $(\nu_{n})$ of the hidden state instead of $(x_{n})$. However, this is not useful in our setting as we focus on observational noise. As we will establish in the next section, when the noise in the observation equation is low, one can consider numerical algorithms under the noise parametrization and ultimately, as the size of the noise goes to 0, recover an algorithm associated to the degenerate noise filtering problem. For this latter problem, we obtain in the subsection that follows expressions analogous to the ones in the current subsection.
@@ 2.3.2 Degenerate Noise
We now consider degenerate observation noise and show the steps for obtaining expression for the sequence of filtering densities w.r.t. appropriate reference measures. We write the SSM under consideration as:
$Y_{n}$ $=H_{n}^{\star}(X_{n}),$
$X_{n}$ $=F_{n}(X_{n-1},\nu_{n}),$
@@
where: $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}+d_{x}}$, $(d_{x},d_{y})\in\mathbb{N}^{2}$, $d_{y}<d_{x}$; $\nu_{n}\in\mathbb{R}^{d_{x}}$ is a sequence of independent noises (across $n$) with positive and smooth Lebesgue density; and $H_{n}^{\star}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$, $F_{n}:\mathbb{R}^{2d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ are sequences of continuously differentiable functions whose Jacobian is of full row rank. Note that we require $d_{y}<d_{x}$ as we aim to work with an embedded manifold.
As above, we set $c_{n}(x_{n})=y_{n}-H_{n}^{\star}(x_{n})$ and define the manifold:
$\mathsf{M}_{n}^{\star}=\{x_{n}\in\mathbb{R}^{d_{x}}:c_{n}(x_{n})=0\}.$ (2.6)
@@
We assume that $\mathbb{R}^{d_{x}}$ is equipped with a metric tensor with a fixed positive definite matrix representation, denoted $M$. We set:
$G_{n}(x_{n})=\textrm{det}(\partial c_{n}(x_{n})M^{-1}\partial c_{n}(x_{n})^{\top})^{-1/2}.$ (2.7)
As before the filter densities w.r.t. $\mu_{\mathsf{M}_{n}^{\star}}$ can be written down as
$\pi_{1}^{\star}(x_{1})\equiv\frac{d\pi_{1}^{\star}}{d\mu_{\mathsf{M}_{1}^{\star}}}(x_{1})=\frac{G_{1}(x_{1})f_{1}(x_{1}|x_{0})}{\int_{\mathsf{M}_{1}^{\star}}G_{1}(x_{1})f_{1}(x_{1}|x_{0})\mu_{\mathsf{M}_{1}^{\star}}(dx_{1})}.$
At any subsequent time $n\geq 2$, we have:
$\pi_{n}^{\star}(x_{n})=\frac{d\pi_{n}^{\star}}{d\mu_{\mathsf{M}_{n}^{\star}}}(x_{n})=\frac{G_{n}(x_{n})\int_{\mathsf{M}_{n-1}^{\star}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}^{\star}(x_{n-1})\mu_{\mathsf{M}_{n-1}^{\star}}(dx_{n-1})}{\int_{\mathsf{M}_{n}^{\star}}G_{n}(x_{n})\int_{\mathsf{M}_{n-1}^{\star}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}^{\star}(x_{n-1})\mu_{\mathsf{M}_{n-1}^{\star}}(dx_{n-1})\mu_{\mathsf{M}_{n}^{\star}}(dx_{n})}.$
Similarly for smoothing density we have
${\pi}_{n}^{\star}(x_{0:n})\equiv\frac{d\widetilde{\pi}_{n}}{d\mu_{\bigotimes_{k=1}^{n}\mathsf{M}^{\star}_{k}}}(x_{0:n})\propto\prod_{k=1}^{n}{G}_{k}(x_{k})f_{k}(x_{k}|x_{k-1}).$ (2.8)
3 Filtering with Linear Observation and Additive Noise
@@
In this section we consider the case of general additive noise (Section 3.1) and derive PFs in the low and degenerate noise cases. We then examine (Section 3.2) the case where the hidden process also involves additive Gaussian noise, similarly to the observation process, and derive in such a setting the optimal proposals for the low and degenerate noise models.
3.1 Reparameterizations and Particle Filters
We consider the SSM given as follows, for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n},$ (3.1)
@@ 3.1 Reparameterizations and Particle Filters
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, $\epsilon_{n}\in\mathbb{R}^{d_{y}}$ forms a sequence of independent noises with positive Lebesgue density $p_{n}(\cdot)$ and $\Delta\geq 0$. For the hidden dynamics we do not impose any strong constraints beyond the existence of a positive Lebesgue transition density which can be evaluated pointwise; it is written as $f_{n}(x_{n}|x_{n-1})$ as before.
@@
In this article we present noise degeneracy as the case that in the observation equation, as $\Delta\downarrow 0$ there is no noise at all. However, one could consider, say $Y_{n}\in\mathbb{R}^{3}$, and an observation equation of the type
$Y_{n}=A_{n}X_{n}+\mathtt{P}(\Delta)\epsilon_{n},$
@@
In this case one could send $\Delta\downarrow 0$ which corresponds to removing the observation noise in the third observation component. This type of scenario could be dealt with in the forthcoming commentary. This is in the sense that one would define a filter (noise parameterization) and a corresponding algorithm in the low noise case and the case $\Delta=0$ and that there is some convergence of the low-noise case to this limiting and degenerate case. The main changes would be mainly notational and hence, for ease of exposition, we keep the simple scenario that we have started with.
Recall our notation for the manifolds:
$\widetilde{\mathsf{M}}_{n}$ $=\{(x_{n},\epsilon_{n})\in\mathbb{R}^{d_{x}+d_{y}}:y_{n}-A_{n}x_{n}-\Delta^{1/2}\epsilon_{n}=0\};$
$\mathsf{M}_{n}^{\star}$ $=\{x_{n}\in\mathbb{R}^{d_{x}}:y_{n}-A_{n}x_{n}=0\}.$
The filtering problems for the low and degenerate noise case will require sampling from $\widetilde{\mathsf{M}}_{n}$ and $\mathsf{M}_{n}^{\star}$ respectively. We will connect the two problems by presenting how one can view $\widetilde{\mathsf{M}}_{n}$ as an appropriate lifting of $\mathsf{M}_{n}^{\star}$. We then discuss how this leads to generic sampling on these manifolds and then implementable PFs. We start from the degenerate case and let $x_{n}^{\star}\in\mathsf{M}_{n}^{\star}$ and let $V_{n}$ denote a $d_{x}\times(d_{x}-d_{y})$ matrix whose columns form an orthonormal basis for the linear subspace $\text{ker}(A_{n})=\{x_{n}\in\mathbb{R}^{d_{x}}:A_{n}x_{n}=0\}$. Thus, we have that:
$\mathsf{M}_{n}^{\star}=x_{n}^{\star}+\mathscr{L}V_{n},$
with $\mathscr{L}$ denoting linear span. The point to note here is that sampling on $\mathsf{M}_{n}^{\star}$ can performed by centering on any $x_{n}^{\star}\in\mathsf{M}_{n}^{\star}$ and adding a random sample defined on $\mathscr{L}V_{n}$. Due to the linear constraint from the observation, our state lives in a lower dimensional manifold, so one can proceed by appropriately sampling $z_{n}\in\mathbb{R}^{d_{x}-d_{y}}$ and then return the linear map $u(z_{n})=x_{n}^{\star}+V_{n}z_{n}$ as a sample in $\mathsf{M}_{n}^{\star}$.
We ‘uplift’ the manifold $\mathsf{M}_{n}^{\star}$ to obtain $\widetilde{\mathsf{M}}_{n}$ by working as follows. Consider starting from the case without noise present ($\varepsilon_{n}=0$) and augment $x_{n}^{\star}$ by setting $(\widetilde{x}_{n}^{\Delta})^{\top}=((x_{n}^{\star})^{\top},0^{\top})$, which is a $(d_{x}+d_{y})$-dimensional vector that obeys the defining constraint of $\widetilde{\mathsf{M}}_{n}$, i.e. $y_{n}=[A_{n},\Delta^{1/2}\mathtt{I}_{d_{y}}]\widetilde{x}_{n}^{\Delta}$. Let $V_{n}^{\Delta}$ be a $(d_{x}+d_{y})\times d_{x}$ matrix whose columns are an orthonormal basis of $\text{ker}([A_{n},\Delta^{1/2}\mathtt{I}_{d_{y}}])$ so that we can write
$\widetilde{\mathsf{M}}_{n}=\widetilde{x}_{n}^{\Delta,0}+\mathscr{L}\big\{V_{n}^{\Delta}\big\}.$
We will set
$V_{n}^{\Delta}=\begin{pmatrix}V_{n}&\Delta^{1/2}A_{n}^{\top}\\ 0&-A_{n}A_{n}^{\top}\end{pmatrix}$ (3.2)
and refer the reader for more details to the Appendix. From a practical point of view at each time $n$ one needs to sample a signal and observation component that lie on the manifold $\widetilde{\mathsf{M}}_{n}$. The first step is to compute a point $x_{n}^{\star}$ solving $A_{n}x_{n}^{\star,0}=y_{n}$ and thus obtain $\widetilde{x}_{n}^{\Delta}$ by augmenting with 0. Then one should sample appropriately $\widetilde{z}\in\mathbb{R}^{d_{x}}$ and use the following mapping:
$u_{n}^{\Delta}(\widetilde{z})=\widetilde{x}_{n}^{\Delta}+V_{n}^{\Delta}\widetilde{z}$, (3.3)
[…]
In both the low noise and degenerate filtering we are primarily interested for the state variables and given this discussion it should be clear one can equivalently define the low noise filtering problem w.r.t a random variable $\widetilde{Z}_{n}\in\mathbb{R}^{d_{x}}$. In order to better connect the two problems we will also use throughout the following notation convention: $\widetilde{z}_{n}^{\top}=(z_{n}^{\top},\overline{z}_{n}^{\top})^{\top}$, with $(z_{n},\overline{z}_{n})\in\mathbb{R}^{d_{x}-d_{y}}\times\mathbb{R}^{d_{y}}$, recalling that $z_{n}$ is the random variable of interest for the degenerate case.
3.1.1 The Low Noise Case
PFs are defined to target a sequence of smoothing distributions, so using (3.3) we will rewrite (2.5) as a function of $\widetilde{z}_{1:n}$ . Note that due to the manifolds considered in this section being linear, the Gram matrices in (2.4), (2.7) are constant and we can write densities w.r.t. Lebesgue measure. Given this construction, at any time $n\geq 1$ we can write the joint smoothing density of $\widetilde{z}_{1:n}\in\mathbb{R}^{nd_{x}}$ w.r.t. $\bigotimes_{i=1}^{n}\lambda_{d_{x}}(d\widetilde{z}_{i})$ as